# CoCo: A Coupled Contrastive Framework for Unsupervised Domain Adaptive Graph Classification

### v3

@@ 1 Introduction
Recently, graph-structured data has flourished in a number of fields including chemistry and bioinformatics. Among various graph-based machine learning tasks, graph classification seeks to predict the properties of whole graphs (Wang et al., 2021; Kong et al., 2022; Ju et al., 2024), and a variety of machine learning algorithms have been put forward for the problem (Yin et al., 2024a, b; Feng et al., 2022; Zhang et al., 2021; Yang et al., 2022). These methods mostly fall under the category of graph neural networks (GNNs). Following the paradigm of message passing (Kipf & Welling, 2017), GNNs learn representations in the graph by stacking multiple neural network layers, each of which transfers semantic information from topological neighbors to centroid nodes. A readout function eventually combines all of the node representations into a graph-level representation. In this way, GNNs are capable of integrating graph structural information into graph representations implicitly, thus facilitating downstream graph classification.
In spite of their certain progress, modern GNN algorithms are typically trained under supervision, necessitating a large quantity of labeled data (Kipf & Welling, 2017; Xu et al., 2019; Bodnar et al., 2021; Baek et al., 2021; Yin & Luo, 2022). However, label annotation in the graph domain is either prohibitively expensive or even impossible to acquire (Xu et al., 2021; Suresh et al., 2021). For instance, ascertaining the pharmacological effect of drug molecule graphs involve costly experiments on living animals. Due to the scarcity of labeled annotations, the bulk of current algorithms performs poorly in practice. To solve this issue, we observe that there are often a substantial number of graph samples from a different but relevant domain and their labels are easily available. In this spirit, this work studies the problem of unsupervised domain adaptive graph classification, a practical task to predict graph properties with both labeled source graphs and unlabeled target graphs.
@@ 2.1 Graph Classification
Graph classification has been a long-standing problem with various applications in social analysis (Yin et al., 2023b, 2022a; Liao et al., 2021; Wu et al., 2020b; Ju et al., 2023a, 2022) and molecular property prediction (Hassani, 2022; Yin et al., 2023a). Early efforts to this problem almost turn to graph kernels including Weisfeiler-Lehman kernel (Shervashidze et al., 2011) and random walk kernel (Kang et al., 2012), which can identity graph substructures through graph decomposition. However, these methods cannot be well applied to large-scale graphs due to high computational costs. To tackle this, a variety of graph neural networks have been put forward in recent years (Kipf & Welling, 2017; Xu et al., 2019; Bodnar et al., 2021; Baek et al., 2021; Ju et al., 2023b). Typically, these algorithms obey the message passing paradigm to iteratively update node representations, followed by a graph pooling function that generates graph-level representations for downstream classification. Hence, these methods usually explore topological information merely in an implicit way. However, recent research has demonstrated that this paradigm is inadequate for detecting structural motifs in graphs such as rings (Long et al., 2021; Chen et al., 2020a; Cosmo et al., 2021). To tackle this issue, besides implicitly exploring graph topological knowledge using the graph convolutional network branch, CoCo utilizes a hierarchical graph kernel network to explicitly explore graph topology, which enhances the classification performance in a domain adaptive framework.
@@ 2.2 Unsupervised Domain Adaptation
The purpose of unsupervised domain adaptation (UDA) is to transfer a model from a label-rich source domain to a label-scarce target domain (He et al., 2022; Huang et al., 2022; Nguyen et al., 2022; Wang et al., 2022). This problem is substantially investigated in the field of computer vision with applications in image classification (Mirza et al., 2022), image retrieval (Zheng et al., 2021) and semantic segmentation (Kundu et al., 2022). In general, domain alignment is the key to successful domain adaptation. The early attempts usually explicitly reduce domain discrepancy using statistical metrics including maximum mean discrepancy (Long et al., 2015) and Wasserstein distance (Shen et al., 2018). Recently, adversarial learning-based approaches are the predominant paradigm for UDA (Long et al., 2018). These methods usually employ a gradient reversal layer (Ganin et al., 2016) to render deep features resistant to domain shifts. In addition, discrimination learning under label scarcity is an important challenge for UDA (Xiao & Zhang, 2021; Liang et al., 2020) and pseudo-labeling techniques are popular strategies (Lee et al., 2013; Assran et al., 2021) to tackle this problem in the target domain. Despite the significant advance of UDA in computer vision, it is still underexplored in the graph domain. In this work, we study the emerging and practical problem of unsupervised domain adaptive graph classification, which employs labeled source graphs to improve the classification performance on unlabeled target graphs. Different from existing domain adaptation methods (Yin et al., 2022b; Mirza et al., 2022; Wei et al., 2021a; Yin et al., 2024c; Pang et al., 2023), our CoCo explores semantics information from different views and conduct coupled contrastive learning for effective domain adaptation.
@@ 3.4 Multi-view Contrastive Learning Framework
where $\tau$ represents a temperature parameter set to $0.5$ as in previous works (He et al., 2020). Through providing challenging views by mining topological information in complementary manners, our proposed CoCo is a special kind of graph contrastive learning (You et al., 2020, 2021; Yin et al., 2022c), hence producing discriminative graph-level representation with sufficient topological information.

### v4

@@ 1 Introduction
Recently, graph-structured data has flourished in a number of fields including chemistry and bioinformatics. Among various graph-based machine learning tasks, graph classification seeks to predict the properties of whole graphs (Wang et al., 2021; Kong et al., 2022), and a variety of machine learning algorithms have been put forward for the problem (Yoo et al., 2022; Feng et al., 2022; Zhang et al., 2021; Yang et al., 2022). These methods mostly fall under the category of graph neural networks (GNNs). Following the paradigm of message passing (Kipf & Welling, 2017), GNNs learn representations in the graph by stacking multiple neural network layers, each of which transfers semantic information from topological neighbors to centroid nodes. A readout function eventually combines all of the node representations into a graph-level representation. In this way, GNNs are capable of integrating graph structural information into graph representations implicitly, thus facilitating downstream graph classification.
In spite of their certain progress, modern GNN algorithms are typically trained under supervision, necessitating a large quantity of labeled data (Kipf & Welling, 2017; Xu et al., 2019; Bodnar et al., 2021; Baek et al., 2021). However, label annotation in the graph domain is either prohibitively expensive or even impossible to acquire (Xu et al., 2021; Suresh et al., 2021). For instance, ascertaining the pharmacological effect of drug molecule graphs involve costly experiments on living animals. Due to the scarcity of labeled annotations, the bulk of current algorithms performs poorly in practice. To solve this issue, we observe that there are often a substantial number of graph samples from a different but relevant domain and their labels are easily available. In this spirit, this work studies the problem of unsupervised domain adaptive graph classification, a practical task to predict graph properties with both labeled source graphs and unlabeled target graphs.
@@ 2.1 Graph Classification
Graph classification has been a long-standing problem with various applications in social analysis (Fan et al., 2019; Song et al., 2016; Liao et al., 2021; Wu et al., 2020b; Ju et al., 2023a, 2022) and molecular property prediction (Hassani, 2022; Yin et al., 2022b). Early efforts to this problem almost turn to graph kernels including Weisfeiler-Lehman kernel (Shervashidze et al., 2011) and random walk kernel (Kang et al., 2012), which can identity graph substructures through graph decomposition. However, these methods cannot be well applied to large-scale graphs due to high computational costs. To tackle this, a variety of graph neural networks have been put forward in recent years (Kipf & Welling, 2017; Xu et al., 2019; Bodnar et al., 2021; Baek et al., 2021; Ju et al., 2023b). Typically, these algorithms obey the message passing paradigm to iteratively update node representations, followed by a graph pooling function that generates graph-level representations for downstream classification. Hence, these methods usually explore topological information merely in an implicit way. However, recent research has demonstrated that this paradigm is inadequate for detecting structural motifs in graphs such as rings (Long et al., 2021; Chen et al., 2020a; Cosmo et al., 2021). To tackle this issue, besides implicitly exploring graph topological knowledge using the graph convolutional network branch, CoCo utilizes a hierarchical graph kernel network to explicitly explore graph topology, which enhances the classification performance in a domain adaptive framework.
@@ 2.2 Unsupervised Domain Adaptation
The purpose of unsupervised domain adaptation (UDA) is to transfer a model from a label-rich source domain to a label-scarce target domain (He et al., 2022; Huang et al., 2022; Nguyen et al., 2022; Wang et al., 2022). This problem is substantially investigated in the field of computer vision with applications in image classification (Mirza et al., 2022), image retrieval (Zheng et al., 2021) and semantic segmentation (Kundu et al., 2022). In general, domain alignment is the key to successful domain adaptation. The early attempts usually explicitly reduce domain discrepancy using statistical metrics including maximum mean discrepancy (Long et al., 2015) and Wasserstein distance (Shen et al., 2018). Recently, adversarial learning-based approaches are the predominant paradigm for UDA (Long et al., 2018). These methods usually employ a gradient reversal layer (Ganin et al., 2016) to render deep features resistant to domain shifts. In addition, discrimination learning under label scarcity is an important challenge for UDA (Xiao & Zhang, 2021; Liang et al., 2020) and pseudo-labeling techniques are popular strategies (Lee et al., 2013; Assran et al., 2021) to tackle this problem in the target domain. Despite the significant advance of UDA in computer vision, it is still underexplored in the graph domain. In this work, we study the emerging and practical problem of unsupervised domain adaptive graph classification, which employs labeled source graphs to improve the classification performance on unlabeled target graphs. Different from existing domain adaptation methods (Yin et al., 2022a; Mirza et al., 2022; Wei et al., 2021a), our CoCo explores semantics information from different views and conduct coupled contrastive learning for effective domain adaptation.
@@ 3.4 Multi-view Contrastive Learning Framework
where $\tau$ represents a temperature parameter set to $0.5$ as in previous works (He et al., 2020). Through providing challenging views by mining topological information in complementary manners, our proposed CoCo is a special kind of graph contrastive learning (You et al., 2020, 2021; Yin et al., 2022b), hence producing discriminative graph-level representation with sufficient topological information.